Question 1: Find a unit vector as below and prove your answer. w=(-5,2,1)
Answer #1

ū= <-5, 2, 1). .:Will = √²+2²+1 25 +4+1 W <-5, 2, 1) 11 W 11 30 =x-Yo Yo Tom unt vector =<- So. 2 / 5 /> Lel. ū. že to .: 11